### v1

# Adaptive Interactive Navigation of Quadruped Robots using Large Language Models

Abstract
Robotic navigation in complex environments remains a critical research challenge. Traditional navigation methods focus on optimal trajectory generation within free space, struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses. To address this gap, we propose an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to reach originally unavailable goals. Specifically, we present a primitive tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences. To ensure robust subtask execution, we adopt reinforcement learning to pre-train a comprehensive skill library containing versatile locomotion and interaction behaviors for motion planning. Furthermore, we introduce an adaptive replanning method featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment. Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan modification in unknown environments. Comprehensive simulations and experiments have demonstrated our method’s effectiveness and adaptivity in diverse scenarios. The supplementary video is available at page: https://youtu.be/W5ttPnSap2g.
Index Terms:
Interactive navigation, large language model, task and motion planning, quadruped robot.
I Introduction
Navigation in diverse and complex environments is a pivotal challenge in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning. Recent research has demonstrated the potential of robotic platforms to traverse cluttered environments and navigate to the desired position, such as utilizing quadrupedal robots [1, 2] for maneuvering across various terrains. Despite the progresses, conventional navigation strategies primarily focus on collision-free trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where collision-free paths may not exist, and the robot must interact with the environment to create viable routes.
To address these limitations, previous works have explored navigation among movable obstacles (NAMO), where the robot can actively interact with and reposition movable obstacles to create viable paths. This NAMO research expands the scope of robotic navigation, enabling operation in highly cluttered and confined environments. However, some prior research [3] focused on global planning within movable obstacles but idealized interactive behavior in simplified simulations, which limited their applicability in real-world situations. To address this limitation, Simon et al. [4] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments. However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse higher hurdles impassable originally, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
Figure 1: Interactive navigation in challenging environments. In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, our method (AINav) can push aside obstacles to create pathways within congested obstacles or utilize tools to traverse excessively high hurdles, tackling interactive navigation tasks in challenging environments.
Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have utilized LLMs combined with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [5, 6]. They formulate the complex interactive tasks as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning considers how to execute the high-level task by a controller to accomplish the task. However, these methods require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations. This limitation restricts their ability to adapt to dynamic and partially observable environments. Closed-loop task planning can respond to new observations through iterative replanning [7, 8]. Nevertheless, previous works commonly focus on passive failure recovery, lacking the ability to proactively utilize new information and adjust plans, which represents a critical aspect for effective navigation in partially observable environments. The primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
To overcome these limitations and challenges, we propose the Adaptive Interactive Navigation (AINav) approach to tackle interactive navigation tasks (Fig. 1). The proposed system demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations in unknown environments. This dual functionality enables effective navigation in unfamiliar and interactive environments using egocentric sensing information, without relying on global environmental knowledge. While recent advancements in robot parkour have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems. The main contributions can be summarized as follows:
  • We propose a primitive skill tree for task planning using LLM, which represents task decomposition as a tree structure with atomic skills. This design enables robust reasoning by exploration of multiple potential solutions and rapid adaptation to new information by tree structure modification.
  • We develop a comprehensive skill library with reinforcement learning (RL) to equip the robot with advanced locomotion and interaction capabilities. Integrated with customized reward design and task curriculum, robots can learn to navigate and manipulate robustly in complex environments.
  • We introduce an adaptive replanning mechanism including two LLM-based modules, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, while the arborist automatically reconstructs the decision tree for optimal task execution.
We validate our method across comprehensive simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments. Through this novel framework, we aim to extend the boundaries of robotic navigation, offering an effective and computationally efficient solution for interactive navigation in complex environments. The remainder of this paper is organized as follows: Section II reviews related work in robotic navigation and interactive planning. Section III presents the proposed AINav framework with specific methods. Section IV describes the experimental setup and results. Finally, Section V concludes the paper and discusses future directions.
II Related Work
II-A Quadruped Robot Navigation
Quadruped robots have demonstrated agile capabilities and terrain traversability in diverse complex environments in recent researches [9, 10]. Based on the superior agility and adaptivity, quadruped robot navigation has made significant progress through various kinds of approaches, including sampling-based method [11], learning-based method [12, 13]. Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [14, 15]. However, current approaches often assume the existence of navigation paths, neglecting the possibility that such paths may not exist in real cluttered scenarios. Zhang et al. [16] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs. However, they aim to adapt to the environment passively rather than actively utilizing interactive objects for task completion. Wu et al. [5] and Xu et al. [6] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively. However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains real-time egocentric observation incrementally.
II-B Navigation Among Movable Obstacles
Navigation among movable obstacles refers to the problem of robot navigation that requires manipulation of the environment. Early research [3] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions. While these approaches are effective in simplified simulation settings, their failure to account for real-world physics significantly restricts their practical applicability. Recent work by Simon et al. [4] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways. This approach incorporates more realistic interaction models, improving applicability in real-world environments. Nevertheless, current NAMO approaches predominantly focus on obstacle removal and do not fully capitalize on the potential of movable obstacles, such as utilizing shorter obstacles as steps to traverse higher barriers. This limitation restricts their ability to address more complex long-horizon tasks in more intricate scenarios.
II-C Closed-Loop Task Planning with LLMs
Due to superior performance in language understanding and commonsense reasoning, LLMs are widely adopted to efficiently solve closed-loop task planning, which refers to generating a series of intermediate steps to achieve the specific goal while modifying the plan based on new observations. Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [17, 18]. However, these methods typically generate a single plan for execution, which often contains errors or infeasible decomposition due to model uncertainties. This increases the need for replanning and hinders computationally efficient planning. Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [19, 7, 20]. However, these methods either lack a replanning mechanism or trigger replanning solely upon encountering failures and concentrate primarily on error correction, overlooking a comprehensive understanding and effective utilization of interactive environments.
Figure 2: An overview of AINav approach, a hierarchical system for interactive navigation. The task planning module processes visual input to generate a subtask skeleton for execution by the motion planning module, integrated with an adaptive replanning mechanism that enables flexible replan triggering and rapid adjustments to the task plan.
III Adaptive Interactive Navigation
Our framework consists of three primary components: LLM-based task planning, motion planning with pre-trained skills, and adaptive replanning, as shown in Fig. 2. The robot first captures the first-person perspective images and extracts scene understanding and object descriptions using a VLM and object detector, respectively. By combining user instructions with environmental information as inputs, the task planning module uses the LLM to construct a primitive tree for reasonable task decomposition and convenient adjustments based on new information, enabling the robot to obtain the optimal high-level primitive skeleton. Leveraging a pre-trained skill library consisting of robust locomotion and interaction skills with RL, the motion planning module executes specific low-level actions following the skeleton to complete the task. Finally, a novel adaptive replanning mechanism interprets new observations from an egocentric perspective to determine whether replanning is necessary, allowing for swift reaction to incremental environmental observations.
III-A Task Planning with Primitive Skill Tree
We propose the primitive skill tree for task planning. The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction, and an evaluator for node evaluation and skeleton selection.
Proposer: object-level to skill-level. Given the user instructions $I$ and environmental information $O$, we first utilize the LLM to propose multiple potential object-level planning strategies $P_{obj}$, where each plan specifies which objects to use and how to leverage them to achieve the task
$\begin{split}P_{obj}=\text{LLM}(I,O).\end{split}$ (1)
Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$.
$\begin{split}P_{skill}=\text{LLM}(P_{obj},S).\end{split}$ (2)
After skill-level plans are generated, we construct the primitive skill tree by merging steps of plans in $P_{skill}$ with identical historical traces. This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected errors, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties. Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object utilization to specific skill planning tends to propose more reasonable and feasible object interaction strategies.
Evaluator: short-term to long-term. To evaluate the proposed plans, we utilize another LLM as an evaluator to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree, which considers the difficulty of primitive skills and the contribution to the specific task. For example, if the robot interacts with objects like pushing movable obstacles to create a free path for navigation, this skill is essential to completing the task, but the interaction is also difficult to complete. If one skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan. If a trajectory in the primitive tree reaches the target point at the end, an additional bonus is given to the leaf node as the terminal reward. Finally, we apply a backup method to update the long-term cumulative reward $Q(n)$ of nodes with
$\begin{split}Q(n)=r(n)+\gamma\cdot\frac{1}{|\text{children}(n)|}\sum_{n_{c}\in\text{children}(n)}Q(n_{c}),\end{split}$ (3)
where $\gamma$ is the discounted factor. By iteratively selecting the child nodes with maximal cumulative reward, we derive the optimal high-level primitive skeleton. This approach also allows for the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be specifically addressed in Section III-C.
III-B Skill Pre-training with Reinforcement Learning
To robustly execute the decomposed tasks, we train a primitive skill library using RL for motion planning. The skill library includes both low-level locomotion skills, such as walking and climbing, and high-level skills like navigation and object-pushing strategies. We first train locomotion policies to develop fundamental movement capabilities on diverse terrains. Building upon the locomotion policies, we employ a hierarchical RL framework to train more advanced strategies, such as navigation policy for collision-free movement in cluttered environments and pushing policy for moving objects to designated positions.
In low-level locomotion skill training, the walking policy training is formulated as a velocity-tracking task, which encourages robots to follow the linear and angular velocity command while adapting to arbitrary terrain. As for the climbing skill, to ensure the robot learns an effective policy, we refine the reward design to encourage the robot to reach the goal while aligning its movement direction toward the goal. Specifically, we add a customized head collision penalty to prevent reliance on the head when climbing up high hurdles, encouraging the robot to use its legs as support instead of the head to learn a more natural and safe climbing action. Additionally, curriculum learning is employed to gradually increase terrain difficulty or hurdle height, enabling the development of robust locomotion policies.
With low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow. Specifically, the pushing skill aims to push boxes to the target pose, and the navigation skill navigates the robot to the target pose while avoiding obstacles. We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill learning, facilitating robust high-level skill learning. All these skills are trained using PPO [21] in the IsaacLab simulation environment. After training these primitive skills, we encapsulate them into APIs for task planning, enabling sequential skill execution to accomplish complex tasks.
III-C Adaptive Replanning with Observation Interpretation
Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot continuously acquires new observations with an egocentric perspective. For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan based on new environmental comprehension. This mechanism consists of two LLM-based roles: an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
The advisor analyzes new environmental information from visual perception and the current plan to determine whether replanning is necessary based on these comprehension updates. We let the advisor conduct an environmental interpretation in three aspects, including failure, new objects, and revaluation. Failure indicates that the current plan has encountered some errors, new objects represent the discovery of a new interactive object that may be useful, and revaluation means reassessing the plan based on updated information about the environment. Unlike previous works that only replan based on failure, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment. For example, when you discover a new object, you need to analyze its geometric and semantic information and decide to utilize it if it is more helpful to your task. As for revaluation, the robot’s understanding of the object information under observation uncertainties will be updated, which will affect the evaluation of the current plan. Based on these three types of comprehension, the advisor will decide whether to replan and replan suggestions.
When the advisor determines that replanning is necessary, it generates specific suggestions for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations. Node addition incorporates new object information, while node pruning removes failed paths. This tree-based architecture enables computationally efficient replanning through its inherent structural flexibility, facilitating real-time task execution. Following structural modifications, the system performs backup operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills. Through the adaptive replanning approach, the robot can respond to new environmental information flexibly and adjust the planning strategies accordingly. By analyzing new observations, the robot can assess their impact on task completion and decide whether adjustments to the current plan are required. This process helps avoid the computational cost of unnecessary replanning.
IV Simulation
In this section, we present the simulation experiments conducted to validate the effectiveness of the proposed method. The experiments are structured into four main parts: (1) the experimental setup, which outlines the simulation environment and interactive navigation tasks; (2) the baseline methods compared to AINav to establish a performance benchmark; (3) the simulation results and analysis; (4) ablation studies to investigate the contribution of individual components of the proposed method.
IV-A Experimental Setup
Simulation Settings. To validate the effectiveness of our proposed method, we conduct extensive simulations in IsaacLab. The simulations are conducted in a $8$m $\times$ $8$m space with fixed obstacles and movable boxes. We design diverse simulation environments corresponding to four interactive navigation tasks, which will be detailed in the following section. We use the Unitree Go2 as the robotic platform to perform the simulations. All experiments are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
Task Description. We design four interactive navigation tasks: box obstruction, box usage, stair building, and an integrated task, as illustrated in Fig. 3. The box obstruction task involves environments where boxes block the path. Box usage and stair building require the robot to interact with boxes to complete navigation tasks. The integrated task combines challenges from the previous three tasks. Each scenario is divided into three levels of difficulty: low, medium, and high. The low-difficulty scenario can be regarded as the traditional navigation task where the goal point can be reached without interaction. Medium difficulty requires an interactive environment, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with immovable boxes, designed to evaluate the replanning performance and efficiency of various methods. The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters. We set the maximum simulation time to $120$s and conducted $10$ repeated experiments for each scenario.
Evaluation Metrics. The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), and execution time (ET) to comprehensively assess the performance of different methods. SR refers to the probability of the robot successfully reaching the target point across 10 trials. OT denotes the total simulation time across all scenarios. If a trial fails, OT is recorded as the maximum simulation time, which is $120$s. OTS represents the overall completion time for successful task completion. PT and ET indicate the time spent on task planning and skill execution under successful completion, respectively. SR, OT, and OTS indicate the comprehensive effectiveness and efficiency of methods, while planning time and execution time reflect the computational efficiency and quality of task planning.
Figure 3: Interactive tasks in simulations.
IV-B Baseline Methods
We compare the performance of the proposed method to the following baselines. Art-planner [11] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks. Hierarchical RL trains a high-level policy to function as the task planning module, which decides which low-level skill to use and the corresponding parameters. The low-level skills utilized in hierarchical RL are identical to those in our method. The reward is inversely related to the distance between the robot and the goal, encouraging the robot to move toward the target. Furthermore, we also select two LLM-based method for comparison. The first is RoboTool [6], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges. Since this method uses open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output. The second method is Tree-planner [7], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems. Because its plan sampling module requires global information, we include the initial object information as input. We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks. For a fair comparison, all methods use the same set of low-level skills as our method.
IV-C Results and Analysis
Simulation results are displayed in Table I. We can observe that our method exhibits higher success rates and less overall time compared to other baselines. The Art-planner achieves high task success rates in low-difficulty scenarios with low execution time, showcasing its effective planning capabilities in traditional navigation tasks. However, in medium-difficulty and high-difficulty scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks. Hierarchical RL can achieve limited success in low-difficulty scenarios of box obstruction, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
TABLE I: Performance Metrics Across Different Scenarios and Methods
Method Metric Box obstruction Box usage Stair building Integrated task
L M H L M H L M H L M H
Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 80% 0% 0%
OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 57.2 120.0 120.0
OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - 41.5 - -
PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - 6.7 - -
ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - 34.8 - -
Hierarchical RL SR 40% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
OT 87.4 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0
OTS 38.5 - - - - - - - - - - -
PT 0.1 - - - - - - - - - - -
ET 38.4 - - - - - - - - - - -
Robotool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 60% 30% 0%
OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 88.0 108.7 120.0
OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 66.6 82.2 -
PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 25.7 -
ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 43.1 56.5 -
Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 70% 30% 0%
OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 90.3 116.6 120.0
OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 77.6 108.5 -
PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 29.3 36.2 -
ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 48.3 72.3 -
Ours SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 80% 60% 50%
OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 69.7 98.9 113.8
OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 57.1 84.9 107.5
PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 20.4 34.1 39.2
ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 36.7 50.8 68.3
  • L = Low difficulty, M = Medium difficulty, H = High difficulty
  • SR = Success Rate, OT = Overall time, PT = Planning Time, ET = Execution Time
Figure 4: Average success rate and overall time across all scenarios of different methods.
Figure 5: Success cases of the proposed method in four interactive navigation tasks. The red phases indicate the robot interacts with objects, while the green phases represent the robot performing replanning based on new environmental observations.
Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, Robotool and tree-planner achieve a higher success rate in interactive navigation tasks compared to traditional navigation methods. However, Robotool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks. While tree-planner leverages a closed-loop design to adjust its decisions based on new observations, its performance is highly dependent on the quality of the sampled plans, resulting in instability in planning results. Specifically, once the tree structure is generated, its inflexibility can hinder the planner from obtaining the optimal solution that does not exist in the original tree or changing to an updated better strategy in response to new observations. Additionally, both methods require omniscient environmental information to generate task decomposition plans, rendering them ineffective in unknown environments. In contrast, our method not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information. This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios, especially in challenging situations that require replanning. Although node evaluation incurs additional planning time, the improved task planning quality and flexible replanning mechanism enable our method to achieve higher success rates and shorter overall completion time compared to other LLM-based approaches. As shown in Fig. 4, AINav demonstrates the highest success rate and the lowest overall completion time across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
We also present successful cases of our method across four tasks in Fig. 5. In the box obstruction scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position. In the box usage and stair building scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which is unreachable for a quadruped robot under normal circumstances. In integrated task scenario, when confronted with various potential interaction tasks, the robot understands the current issue by VLM and intelligently interacts with the appropriate objects to facilitate task completion. Furthermore, we can observe that the robot is capable of responding to updated information, such as when the box fails to move or the robot identifies new objects and tasks. By leveraging an efficient closed-loop system, the proposed approach enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
IV-D Ablation Studies
Figure 6: Comparison results in ablation study.
To further investigate the effectiveness of the proposed method, we conduct ablation study to compare AINav with the following four variants. 1) AINav-Single, which only generates a single plan with LLM in the task planning module; 2) AINav-Skill, which directly generates skills without incorporating object-centric reasoning; 3) AINav-No, which removes the replanning module; 4) AINav-F, which only replans when a failure is encountered. We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building, which need interaction for task completion, to compare the success rate and overall time. The results of ablation studies are presented in Fig. 6. AINav-Single exhibits low success rates, as the single plan generated by the LLM is often flawed or inefficient due to inherent uncertainties of LLM, leading to task failures. Benefiting from the tree structure that enhances the robustness and planning quality, AINav-Skill achieves a reasonable success rate. However, it lacks a comprehensive understanding of object utilization in complex scenarios, leading to task decomposition plans inexecutable. AINav-No performs poorly because it lacks a replanning phase, making it unable to adapt to new observations. Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, often resulting in excessive time spent and task timeouts. By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving higher success rates and improved time efficiency.
IV-E Performance of Primitive Skills
(a) Climbing up/down skill
(b) Navigation skill
(c) Pushing skill: small rectangle box
(d) Pushing skill: big square box
Figure 7: Display of diverse RL-driven skills. (a) and (b) show the climbing and navigation skills adapted to various terrains, while (c) and (d) showcase robust interaction capabilities with boxes of different shapes and sizes.
We visualize the performance of various skills in Fig. 7. In Fig. 7(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks. Fig. 7(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment. In Fig. 7(c) and Fig. 7(d), we showcase the robot pushes boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy. The box-pushing process typically consists of three steps: first, the robot approaches the box and positions itself so that the box is between the robot and the target position, preparing for the subsequent pushing action; second, the robot pushes the box toward the target point; and finally, the robot fine-tunes the box’s position to complete the task. These versatile skills collectively form a comprehensive skill library that enables the robot to handle various sub-tasks in complex navigation scenarios.
V Experiment
Figure 8: Box-pushing policy in different configurations.
(a) Different scenarios for real-world experiments.
(b) Box obstruction
(c) Box utilization: indoor
(d) Box utilization: outdoor
Figure 9: Visualization of real-world experiments. (a) shows various scenarios in real-world experiments. (b), (c) and (d) illustrate case studies for each scenario, in which the top-left subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
We also conduct real-world experiments to validate the performance of AINav in practical situations. We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection. We evaluate the effectiveness of our method across three interactive navigation scenarios, as illustrated in Fig. 9(a). The first scenario involves an indoor obstacle obstruction setting, simulating a navigation task in a narrow corridor blocked by movable obstacles, where the robot needs to remove the box to reach the target. The second scenario, also set in a confined indoor corridor, featured a fixed $0.35$m-height platform obstructing the path. Since we assume the maximum climbing height for the robot is $0.3$ meters, the robot needs to reason how to traverse heights that exceed this capability limit by interacting with objects. The third scenario is box utilization in an outdoor environment, setting with fixed obstacles blocking the path and $0.45$m height, presenting a more difficult challenge with a different layout. Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain accurate object descriptions, thereby mitigating the risk of task failure due to perception errors.
Fig. 9(b) demonstrates the experiment on obstacle obstruction. Initially, the robot detects a blocking box and navigates around it using local point cloud information and its navigation policy to approach the box. Subsequently, a pushing skill is employed to move the obstacle aside, creating a path to the target. Fig. 9(c) illustrates how the robot uses a box to overcome a higher obstacle. After visually detecting the box, the task planning module decides to push the box next to the high platform as a step, enabling the robot to climb over the challenging obstacle and reach the target. Fig. 9(d) shows the robot completing the interactive navigation task in the outdoor scenario, employing task decomposition and skill execution to achieve the target, overcoming challenges that traditional navigation methods could not address. Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction. These above real-world experiments demonstrate our method’s capability to complete navigation tasks in challenging environments. Further details are available in our supplementary video. We conduct $5$ repeated experiments in each scenarios, and the success rates are $100\%$ ($/5$ 5), $80\%$ ($4/5$), and $80\%$ ($4/5$), respectively, highlighting the robustness and superiority of AINav in various interactive tasks and both indoor and outdoor settings.
To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig. 8. We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement. This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
VI Conclusion
We introduce an LLM-based adaptive interactive navigation approach using a quadruped robot, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals. Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning. The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained. Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal. Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations. This instructive framework has the potential to enhance robotic navigation capabilities, offering an effective and computationally efficient solution for interactive navigation in challenging environments. Future directions encompass developing more generalizable interaction skills to accommodate broader tasks and scenarios, and achieving tighter integration between LLM/VLM and versatile control policies for more intelligent embodied planning and control.

### v2

@@
-Adaptive Interactive Navigation of Quadruped Robots using Large Language Models
+AINav: Interactive Navigation via LLM-based Primitive Skill Tree and Adaptive Replanning
 Abstract
 Robotic navigation in complex environments remains a critical research challenge.
-Traditional navigation methods focus on optimal trajectory generation within free space, struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses.
-To address this gap, we propose an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to reach originally unavailable goals.
-Specifically, we present a primitive tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences.
+Traditional navigation methods focus on optimal trajectory generation within fixed free workspace, therefore struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses.
+To address this problem, we propose AINav, an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to achieve originally unreachable goals.
+Specifically, we present a primitive skill tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences.
 To ensure robust subtask execution, we adopt reinforcement learning to pre-train a comprehensive skill library containing versatile locomotion and interaction behaviors for motion planning.
-Furthermore, we introduce an adaptive replanning method featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment.
-Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan modification in unknown environments.
-Comprehensive simulations and experiments have demonstrated our method’s effectiveness and adaptivity in diverse scenarios.
-The supplementary video is available at page: https://youtu.be/W5ttPnSap2g.
-Index Terms:
-Interactive navigation, large language model, task and motion planning, quadruped robot.
+Furthermore, we introduce an adaptive replanning approach featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment.
+Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan adaptation in a priori unknown environments.
+Comprehensive simulations and experiments have demonstrated AINav’s effectiveness and adaptivity in diverse scenarios.
+The supplementary video is available at: https://youtu.be/CjXm5KFx9AI.
 I Introduction
-Navigation in diverse and complex environments is a pivotal challenge in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning.
-Recent research has demonstrated the potential of robotic platforms to traverse cluttered environments and navigate to the desired position, such as utilizing quadrupedal robots [1, 2] for maneuvering across various terrains.
-Despite the progresses, conventional navigation strategies primarily focus on collision-free trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where collision-free paths may not exist, and the robot must interact with the environment to create viable routes.
-To address these limitations, previous works have explored navigation among movable obstacles (NAMO), where the robot can actively interact with and reposition movable obstacles to create viable paths.
-This NAMO research expands the scope of robotic navigation, enabling operation in highly cluttered and confined environments.
-However, some prior research [3] focused on global planning within movable obstacles but idealized interactive behavior in simplified simulations, which limited their applicability in real-world situations.
+Navigation in diverse and complex environments is a pivotal problem in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning.
+Recent research has demonstrated the potential of quadruped robots to traverse cluttered environments and navigate to the desired position on complex terrains [hoeller2024anymal, han2024lifelike].
+Despite the progresses, conventional navigation strategies primarily focus on optimal trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where feasible paths may not exist a priori, and the robot needs to interact with the environment to create viable routes.
+To tackle this challenge, previous works have explored strategies for a special case, namely navigation among movable obstacles (NAMO), where the robot can actively interact with movable obstacles to create viable paths.
+However, prior research [muguira2023visibility] primarily focused on global planning and assumed deterministic object dynamics during interaction in simplified simulations, which limited their applicability in real-world situations.
 To address this limitation, Simon et al.
-[4] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments.
-However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse higher hurdles impassable originally, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
+[armleder2024tactile] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments.
+However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse originally impassable hurdles, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
 Figure 1: Interactive navigation in challenging environments.
-In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, our method (AINav) can push aside obstacles to create pathways within congested obstacles or utilize tools to traverse excessively high hurdles, tackling interactive navigation tasks in challenging environments.
-Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have utilized LLMs combined with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [5, 6].
-They formulate the complex interactive tasks as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning considers how to execute the high-level task by a controller to accomplish the task.
-However, these methods require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations.
+In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, AINav can push aside obstacles to create pathways within congested obstacles or utilize obstacles as tools to traverse excessively high hurdles, enabling interactive navigation in challenging environments.
+Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have combined LLMs with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [ouyang2024long, xu2023creative].
+Specifically, the interactive tasks are formulated as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning designs controllers to execute the high-level tasks.
+Existing TAMP methods usually require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations.
 This limitation restricts their ability to adapt to dynamic and partially observable environments.
-Closed-loop task planning can respond to new observations through iterative replanning [7, 8].
+Closed-loop task planning can respond to new observations through iterative replanning [hu2023tree, wang2023describe].
 Nevertheless, previous works commonly focus on passive failure recovery, lacking the ability to proactively utilize new information and adjust plans, which represents a critical aspect for effective navigation in partially observable environments.
-The primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
-To overcome these limitations and challenges, we propose the Adaptive Interactive Navigation (AINav) approach to tackle interactive navigation tasks (Fig.
-1).
-The proposed system demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations in unknown environments.
+Therefore, the primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
+To overcome these challenges, we propose the Adaptive Interactive Navigation (AINav), a closed-loop system that demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations (Fig.˜1).
 This dual functionality enables effective navigation in unfamiliar and interactive environments using egocentric sensing information, without relying on global environmental knowledge.
-While recent advancements in robot parkour have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems.
+While recent advancements in robot parkour [cheng2024extreme] have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems.
 The main contributions can be summarized as follows:
@@
   • We develop a comprehensive skill library with reinforcement learning (RL) to equip the robot with advanced locomotion and interaction capabilities.
-Integrated with customized reward design and task curriculum, robots can learn to navigate and manipulate robustly in complex environments.
-  • We introduce an adaptive replanning mechanism including two LLM-based modules, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, while the arborist automatically reconstructs the decision tree for optimal task execution.
-We validate our method across comprehensive simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments.
+Integrated with customized reward design and task curriculum, robots can learn to navigate and interact robustly in complex environments.
+  • We introduce an adaptive replanning mechanism including two LLM-based agents, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, and the arborist automatically adjusts the decision tree for optimal task execution.
+We validate AINav across various simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments.
 Through this novel framework, we aim to extend the boundaries of robotic navigation, offering an effective and computationally efficient solution for interactive navigation in complex environments.
 The remainder of this paper is organized as follows: Section II reviews related work in robotic navigation and interactive planning.
-Section III presents the proposed AINav framework with specific methods.
-Section IV describes the experimental setup and results.
-Finally, Section V concludes the paper and discusses future directions.
+Section III introduces the problem formulation of interactive navigation.
+Section IV presents the proposed AINav framework.
+Section V describes the experimental setup and results.
+Finally, Section VI concludes the manuscript and discusses future directions.
 II Related Work
-II-A Quadruped Robot Navigation
-Quadruped robots have demonstrated agile capabilities and terrain traversability in diverse complex environments in recent researches [9, 10].
-Based on the superior agility and adaptivity, quadruped robot navigation has made significant progress through various kinds of approaches, including sampling-based method [11], learning-based method [12, 13].
-Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [14, 15].
+II-A Robot Navigation in Complex Environments
+Navigation in complex environments requires robots to traverse cluttered spaces, dynamically avoid obstacles, and adapt to irregular terrains [hoeller2024anymal].
+Leveraging the superior agility and adaptivity of quadruped robots, significant progress has been made to complete navigation tasks across diverse complex terrain, including sampling-based methods [wellhausen2023artplanner], learning-based methods [rudin2022advanced].
+Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [zhu2024cross, shah2023vint].
 However, current approaches often assume the existence of navigation paths, neglecting the possibility that such paths may not exist in real cluttered scenarios.
 Zhang et al.
-[16] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs.
+[zhang2024interactive] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs.
 However, they aim to adapt to the environment passively rather than actively utilizing interactive objects for task completion.
 Wu et al.
-[5] and Xu et al.
-[6] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively.
-However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains real-time egocentric observation incrementally.
+[ouyang2024long] and Xu et al.
+[xu2023creative] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively.
+However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains egocentric observation incrementally.
 II-B Navigation Among Movable Obstacles
 Navigation among movable obstacles refers to the problem of robot navigation that requires manipulation of the environment.
-Early research [3] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions.
+Early research [muguira2023visibility] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions.
 While these approaches are effective in simplified simulation settings, their failure to account for real-world physics significantly restricts their practical applicability.
 Recent work by Simon et al.
-[4] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways.
+[armleder2024tactile] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways.
 This approach incorporates more realistic interaction models, improving applicability in real-world environments.
@@
 This limitation restricts their ability to address more complex long-horizon tasks in more intricate scenarios.
+Figure 2: An overview of AINav approach, a hierarchical system for interactive navigation.
+The task planning module processes visual input to generate a subtask skeleton for execution by the motion planning module, integrated with an adaptive replanning mechanism that enables flexible replan triggering and rapid adjustments to the task plan.
 II-C Closed-Loop Task Planning with LLMs
 Due to superior performance in language understanding and commonsense reasoning, LLMs are widely adopted to efficiently solve closed-loop task planning, which refers to generating a series of intermediate steps to achieve the specific goal while modifying the plan based on new observations.
-Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [17, 18].
+Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [liu2023reflect, wang2024llm].
 However, these methods typically generate a single plan for execution, which often contains errors or infeasible decomposition due to model uncertainties.
 This increases the need for replanning and hinders computationally efficient planning.
-Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [19, 7, 20].
+Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [yao2024tree, hu2023tree, zhao2024large].
+LLM-BT [zhou2024llm] combines LLM with behavior trees to automatically update the tree, addressing the reactive rigidity of traditional methods to adapt to environmental changes.
 However, these methods either lack a replanning mechanism or trigger replanning solely upon encountering failures and concentrate primarily on error correction, overlooking a comprehensive understanding and effective utilization of interactive environments.
-Figure 2: An overview of AINav approach, a hierarchical system for interactive navigation.
-The task planning module processes visual input to generate a subtask skeleton for execution by the motion planning module, integrated with an adaptive replanning mechanism that enables flexible replan triggering and rapid adjustments to the task plan.
-III Adaptive Interactive Navigation
-Our framework consists of three primary components: LLM-based task planning, motion planning with pre-trained skills, and adaptive replanning, as shown in Fig.
-2.
-The robot first captures the first-person perspective images and extracts scene understanding and object descriptions using a VLM and object detector, respectively.
-By combining user instructions with environmental information as inputs, the task planning module uses the LLM to construct a primitive tree for reasonable task decomposition and convenient adjustments based on new information, enabling the robot to obtain the optimal high-level primitive skeleton.
-Leveraging a pre-trained skill library consisting of robust locomotion and interaction skills with RL, the motion planning module executes specific low-level actions following the skeleton to complete the task.
-Finally, a novel adaptive replanning mechanism interprets new observations from an egocentric perspective to determine whether replanning is necessary, allowing for swift reaction to incremental environmental observations.
-III-A Task Planning with Primitive Skill Tree
-We propose the primitive skill tree for task planning.
-The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction, and an evaluator for node evaluation and skeleton selection.
+Figure 3: Task planning with the proposer and the evaluator.
+III Problem Formulation
+We formalize interactive navigation as a 6-tuple TAMP problem $(\mathcal{X}_{init},\mathcal{X}_{goal},\mathcal{O}_{s},\mathcal{O}_{m},\mathcal{A},\mathcal{C})$, with robot’s initial position $\mathcal{X}_{init}$, navigation goal position $\mathcal{X}_{goal}$, fixed obstacles $\mathcal{O}_{f}$, movable obstacles $\mathcal{O}_{m}$, action space $\mathcal{A}$, and constraints $\mathcal{C}$.
+The objective of interactive navigation is to execute a sequence of actions $\pi=[a_{1},\ldots,a_{n}]$ that enables the robot to move from initial position $\mathcal{X}_{init}$ to goal position $\mathcal{X}_{goal}$, where each action $a_{i}=(p_{i},q_{i})\in\mathcal{A}$ ($i=1,\ldots,n$) consists of a primitive skill $p_{i}$ from the pre-defined skill library and continuous parameters $q_{i}$ for execution policy.
+Finally, each action $a_{i}$ is executed by a pre-trained policy that generates the joint control commands for quadruped robot actuation.
+We assume that the robot knows whether the objects within its field of view are movable, i.e., whether an object belongs to $\mathcal{O}_{f}$ or $\mathcal{O}_{m}$.
+Unlike traditional navigation problems that assume a fixed free space and typically the existence of a feasible path to the goal, interactive navigation tasks consider scenarios where no feasible path exists in the original workspace, and the robot needs to proactively change the workspace through interactions to create feasible paths.
+We propose a hierarchical planning framework, AINav, which decides the primitive skills to execute for task planning and specifies execution details for motion planning, jointly optimizing discrete skill selection and continuous execution parameters to tackle interactive navigation problems.
+IV Adaptive Interactive Navigation
+IV-A Framework Overview
+AINav consists of three primary components: LLM-based task planning, motion planning with pre-trained skill library, and adaptive replanning, as shown in Fig.˜2.
+The robot first captures the first-person perspective images, and extracts scene understanding (e.g., whether pathways are obstructed) using a VLM, and an object detector provides object descriptions including spatial positions and geometric properties.
+By combining user instructions, including navigation goal, skill library, and constraints, with environmental information as inputs, the task planning module uses the LLM to construct a primitive skill tree for reasoning and task decomposition, enabling the robot to obtain the optimal high-level primitive skeleton.
+Taking the height data from lidar as input, the motion planning module executes specific low-level actions following the skeleton by leveraging a pre-trained skill library consisting of robust locomotion and interaction skills.
+Finally, a novel adaptive replanning mechanism interprets new egocentric observations, not only determining whether replanning is necessary but also adjusting the current plan for swift adaptation to incremental environmental observations.
+The detailed prompt of the proposed LLM-based proposer, evaluator, advisor, and arborist can be found in Section˜-A.
+IV-B Task Planning with Primitive Skill Tree
+We propose the primitive skill tree for task planning shown in Fig.˜3.
+The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction denoted as $\text{LLM}_{p}$, and an evaluator for node evaluation and skeleton selection, denoted as $\text{LLM}_{e}$.
 Proposer: object-level to skill-level.
-Given the user instructions $I$ and environmental information $O$, we first utilize the LLM to propose multiple potential object-level planning strategies $P_{obj}$, where each plan specifies which objects to use and how to leverage them to achieve the task
-$\begin{split}P_{obj}=\text{LLM}(I,O).\end{split}$ (1)
-Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$.
-$\begin{split}P_{skill}=\text{LLM}(P_{obj},S).\end{split}$ (2)
+Given the user instructions $I$ and environmental information $O$, we first utilize $\text{LLM}_{p}$ to propose multiple potential object-level strategies $P_{obj}$, where each plan specifies which objects to interact with and how to leverage them to achieve the navigation goal,
+$\begin{split}P_{obj}=\text{LLM}_{p}(I,O).\end{split}$ (1)
+Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$,
+$\begin{split}P_{skill}=\text{LLM}_{p}(P_{obj},S).\end{split}$ (2)
+Each primitive skill consists of the skill name and symbolic parameters, including walk (target position), climb (target position), navigate (target position), and push (object id, target position).
 After skill-level plans are generated, we construct the primitive skill tree by merging steps of plans in $P_{skill}$ with identical historical traces.
-This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected errors, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties.
-Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object utilization to specific skill planning tends to propose more reasonable and feasible object interaction strategies.
+Specifically, each plan path is iteratively added to the tree.
+If a sequence of primitive skills and parameters already exists in the tree, new plans are merged with the existing structure at that common node.
+Otherwise, a new tree branch is created.
+This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected situations, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties.
+Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object reasoning and utilization to specific skill planning tends to propose more reasonable and feasible interaction strategies.
 Evaluator: short-term to long-term.
-To evaluate the proposed plans, we utilize another LLM as an evaluator to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree, which considers the difficulty of primitive skills and the contribution to the specific task.
+To evaluate the proposed plans, we utilize $\text{LLM}_{e}$ to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree generated by proposer $\text{LLM}_{p}$, which considers the contribution to the specific task and the difficulty of pre-defined primitive skills.
 For example, if the robot interacts with objects like pushing movable obstacles to create a free path for navigation, this skill is essential to completing the task, but the interaction is also difficult to complete.
-If one skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan.
-If a trajectory in the primitive tree reaches the target point at the end, an additional bonus is given to the leaf node as the terminal reward.
-Finally, we apply a backup method to update the long-term cumulative reward $Q(n)$ of nodes with
+If a skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan.
+If a trajectory in the primitive tree reaches the navigation goal at the end, an additional bonus is given to the leaf node as the terminal reward.
+Finally, we apply a backpropagation method to update the long-term cumulative reward $Q(n)$ of nodes as follows,
 $\begin{split}Q(n)=r(n)+\gamma\cdot\frac{1}{|\text{children}(n)|}\sum_{n_{c}\in\text{children}(n)}Q(n_{c}),\end{split}$ (3)
-where $\gamma$ is the discounted factor.
+where $\gamma$ is the discount factor.
 By iteratively selecting the child nodes with maximal cumulative reward, we derive the optimal high-level primitive skeleton.
-This approach also allows for the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be specifically addressed in Section III-C.
-III-B Skill Pre-training with Reinforcement Learning
+The symbolic parameters are then converted into precise coordinates using the LLM for execution, which is detailed in Section˜-A.
+This tree-structured approach also allows the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be described in Section˜IV-D.
+IV-C Skill Pre-training with Reinforcement Learning
 To robustly execute the decomposed tasks, we train a primitive skill library using RL for motion planning.
 The skill library includes both low-level locomotion skills, such as walking and climbing, and high-level skills like navigation and object-pushing strategies.
-We first train locomotion policies to develop fundamental movement capabilities on diverse terrains.
-Building upon the locomotion policies, we employ a hierarchical RL framework to train more advanced strategies, such as navigation policy for collision-free movement in cluttered environments and pushing policy for moving objects to designated positions.
 In low-level locomotion skill training, the walking policy training is formulated as a velocity-tracking task, which encourages robots to follow the linear and angular velocity command while adapting to arbitrary terrain.
@@
 Additionally, curriculum learning is employed to gradually increase terrain difficulty or hurdle height, enabling the development of robust locomotion policies.
-With low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow.
+Building upon low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow.
 Specifically, the pushing skill aims to push boxes to the target pose, and the navigation skill navigates the robot to the target pose while avoiding obstacles.
-We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill learning, facilitating robust high-level skill learning.
-All these skills are trained using PPO [21] in the IsaacLab simulation environment.
+We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill acquisition, facilitating robust high-level skill learning.
+All these skills are trained using PPO [schulman2017proximal] in the IsaacLab simulation environment.
+The reward terms of all skills are detailed in Section˜-B.
 After training these primitive skills, we encapsulate them into APIs for task planning, enabling sequential skill execution to accomplish complex tasks.
-III-C Adaptive Replanning with Observation Interpretation
-Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot continuously acquires new observations with an egocentric perspective.
-For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan based on new environmental comprehension.
-This mechanism consists of two LLM-based roles: an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
-The advisor analyzes new environmental information from visual perception and the current plan to determine whether replanning is necessary based on these comprehension updates.
-We let the advisor conduct an environmental interpretation in three aspects, including failure, new objects, and revaluation.
-Failure indicates that the current plan has encountered some errors, new objects represent the discovery of a new interactive object that may be useful, and revaluation means reassessing the plan based on updated information about the environment.
-Unlike previous works that only replan based on failure, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment.
-For example, when you discover a new object, you need to analyze its geometric and semantic information and decide to utilize it if it is more helpful to your task.
-As for revaluation, the robot’s understanding of the object information under observation uncertainties will be updated, which will affect the evaluation of the current plan.
-Based on these three types of comprehension, the advisor will decide whether to replan and replan suggestions.
-When the advisor determines that replanning is necessary, it generates specific suggestions for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations.
-Node addition incorporates new object information, while node pruning removes failed paths.
+IV-D Adaptive Replanning with Observation Interpretation
+Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot acquires new observations with an egocentric perspective.
+For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan when needed.
+This mechanism consists of two LLM-based agents, an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
+Taking new environmental observations and the current plan as inputs, the advisor conducts an observation interpretation and determines whether replanning is necessary.
+We propose three aspects of observation interpretation, including failure, new objects, and revaluation.
+Failure indicates the current plan is practically infeasible, new objects represent the perception of a new object that may be useful, and revaluation means reassessing the plan based on updated environmental information.
+Unlike previous works that only replan when encountering a failure situation, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment.
+For example, upon detecting a new object, the robot needs to assess its geometric properties and determine whether utilizing it would enhance task performance.
+As for revaluation, perceived object properties, such as whether movable, may change during the robot’s movement, leading to updated affordances that may necessitate revaluating the interaction strategies in the current plan.
+Triggered by one of these three interpretations, the advisor decides whether to replan by comparing potential improvements against the current plan.
+If the advisor determines that replanning is necessary, it generates specific suggestions about adjustments to the current plan according to the interpretation of the new observations for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations.
+Node addition takes new objects into consideration in the primitive tree, while node pruning removes failed paths.
 This tree-based architecture enables computationally efficient replanning through its inherent structural flexibility, facilitating real-time task execution.
-Following structural modifications, the system performs backup operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills.
+Following structural modifications, the system performs backpropagation operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills.
 Through the adaptive replanning approach, the robot can respond to new environmental information flexibly and adjust the planning strategies accordingly.
-By analyzing new observations, the robot can assess their impact on task completion and decide whether adjustments to the current plan are required.
-This process helps avoid the computational cost of unnecessary replanning.
-IV Simulation
-In this section, we present the simulation experiments conducted to validate the effectiveness of the proposed method.
-The experiments are structured into four main parts: (1) the experimental setup, which outlines the simulation environment and interactive navigation tasks; (2) the baseline methods compared to AINav to establish a performance benchmark; (3) the simulation results and analysis; (4) ablation studies to investigate the contribution of individual components of the proposed method.
-IV-A Experimental Setup
+V Simulation
+Figure 4: Simulation environments for interactive navigation tasks.
+V-A Experimental Setup
 Simulation Settings.
-To validate the effectiveness of our proposed method, we conduct extensive simulations in IsaacLab.
-The simulations are conducted in a $8$m $\times$ $8$m space with fixed obstacles and movable boxes.
-We design diverse simulation environments corresponding to four interactive navigation tasks, which will be detailed in the following section.
-We use the Unitree Go2 as the robotic platform to perform the simulations.
-All experiments are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
-Task Description.
-We design four interactive navigation tasks: box obstruction, box usage, stair building, and an integrated task, as illustrated in Fig.
-3.
-The box obstruction task involves environments where boxes block the path.
-Box usage and stair building require the robot to interact with boxes to complete navigation tasks.
-The integrated task combines challenges from the previous three tasks.
-Each scenario is divided into three levels of difficulty: low, medium, and high.
-The low-difficulty scenario can be regarded as the traditional navigation task where the goal point can be reached without interaction.
-Medium difficulty requires an interactive environment, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with immovable boxes, designed to evaluate the replanning performance and efficiency of various methods.
-The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters.
-We set the maximum simulation time to $120$s and conducted $10$ repeated experiments for each scenario.
-Evaluation Metrics.
-The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), and execution time (ET) to comprehensively assess the performance of different methods.
-SR refers to the probability of the robot successfully reaching the target point across 10 trials.
-OT denotes the total simulation time across all scenarios.
-If a trial fails, OT is recorded as the maximum simulation time, which is $120$s.
-OTS represents the overall completion time for successful task completion.
-PT and ET indicate the time spent on task planning and skill execution under successful completion, respectively.
-SR, OT, and OTS indicate the comprehensive effectiveness and efficiency of methods, while planning time and execution time reflect the computational efficiency and quality of task planning.
-Figure 3: Interactive tasks in simulations.
-IV-B Baseline Methods
-We compare the performance of the proposed method to the following baselines.
-Art-planner [11] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks.
-Hierarchical RL trains a high-level policy to function as the task planning module, which decides which low-level skill to use and the corresponding parameters.
-The low-level skills utilized in hierarchical RL are identical to those in our method.
-The reward is inversely related to the distance between the robot and the goal, encouraging the robot to move toward the target.
-Furthermore, we also select two LLM-based method for comparison.
-The first is RoboTool [6], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges.
-Since this method uses open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output.
-The second method is Tree-planner [7], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems.
-Because its plan sampling module requires global information, we include the initial object information as input.
-We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks.
-For a fair comparison, all methods use the same set of low-level skills as our method.
-IV-C Results and Analysis
-Simulation results are displayed in Table I.
-We can observe that our method exhibits higher success rates and less overall time compared to other baselines.
-The Art-planner achieves high task success rates in low-difficulty scenarios with low execution time, showcasing its effective planning capabilities in traditional navigation tasks.
-However, in medium-difficulty and high-difficulty scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks.
-Hierarchical RL can achieve limited success in low-difficulty scenarios of box obstruction, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
+We conduct simulations in IsaacSim to validate the effectiveness of AINav.
+The simulation environment includes the maze, indoor Matterport3d dataset, urban ruin, and field ruin, as shown in Fig.˜4.
+In the maze environment, we design diverse tasks and scenarios, which will be detailed in the following section.
+For the other three environments, we test AINav in simulated realistic indoor and outdoor scenarios, including cluttered domestic environments and ruin scenes.
+We use the GPT-4o model as the LLM backbone for all methods in our experiments, the CLIP model as VLM, the YOLOv8 model as the detection model, and the Unitree Go2 as the robotic platform to perform the simulations.
+To simplify visual processing, we assume that the robot can obtain noisy dimensions of objects within the robot’s field of view, modeled as white noise with a standard deviation proportional to the distance between the robot and the objects.
+All simulations are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
 TABLE I: Performance metrics across different scenarios and methods
-Method Metric Box obstruction Box usage Stair building Integrated task
-L M H L M H L M H L M H
-Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 80% 0% 0%
-OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 57.2 120.0 120.0
-OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - 41.5 - -
-PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - 6.7 - -
-ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - 34.8 - -
+Method Metric Box obstruction Box usage Stair building Matterport Urban ruin Field ruin
+L M H L M H L M H
+Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 0% 0% 0%
+OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 120.0 120.0 120.0
+OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - - - -
+PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - - - -
+ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - - - -
+TLS $\downarrow$ 14.8 - - 17.4 - - 16.8 - - - - -
 Hierarchical RL SR 40% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
@@
 ET 38.4 - - - - - - - - - - -
-Robotool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 60% 30% 0%
-OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 88.0 108.7 120.0
-OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 66.6 82.2 -
-PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 25.7 -
-ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 43.1 56.5 -
-Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 70% 30% 0%
-OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 90.3 116.6 120.0
-OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 77.6 108.5 -
-PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 29.3 36.2 -
-ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 48.3 72.3 -
-Ours SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 80% 60% 50%
-OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 69.7 98.9 113.8
-OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 57.1 84.9 107.5
-PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 20.4 34.1 39.2
-ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 36.7 50.8 68.3
-  • L = Low difficulty, M = Medium difficulty, H = High difficulty
-  • SR = Success Rate, OT = Overall time, PT = Planning Time, ET = Execution Time
+TLS 24.1 - - - - - - - - - - -
+RoboTool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 30% 0% 20%
+OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 110.3 120.0 113.9
+OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 87.8 - 89.7
+PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 - 28.4
+ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 64.3 - 61.3
+TLS 19.6 24.3 29.5 19.2 23.7 26.7 20.9 31.4 - 43.4 - 36.7
+Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 50% 0% 0%
+OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 103.6 120.0 120.0
+OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 87.2 - -
+PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 32.2 - -
+ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 55.0 - -
+TLS 19.1 22.8 34.6 16.7 28.5 34.0 22.3 24.9 46.1 38.4 - -
+AINav (Ours) SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 90% 30% 50%
+OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 77.3 113.1 97.8
+OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 72.7 97.1 75.6
+PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 19.5 28.2 30.8
+ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 53.2 68.9 44.8
+TLS 17.4 18.9 25.6 16.2 19.2 26.4 17.1 23.4 31.3 34.9 24.6 21.5
+  • L = Low difficulty, M = Medium difficulty, H = High difficulty, SR = Success rate, OT = Overall time, OTS = Overall time under successful trials, PT = Planning time, ET = Execution time, TLS = Trajectory length under successful trials
+Task Description.
+In the maze environment, we design three interactive navigation tasks: box obstruction, box usage, and stair building, as illustrated in Fig.˜6.
+The box obstruction task involves environments where boxes block the path.
+Box usage and stair building require the robot to utilize boxes as tools to complete navigation tasks.
+Each scenario is divided into three levels of difficulty: low, medium, and high.
+The low-difficulty scenario removes the interaction tasks for simple navigation, which can be regarded as traditional navigation.
+Medium difficulty requires an interaction, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with heavy boxes that are difficult to push, designed to evaluate the replanning performance and efficiency of various methods.
+In the Matterport3d and ruin environments, we designate the start and navigation goal shown in Fig.˜7 to test the effectiveness of AINav in diverse environments.
+The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters as a constraint.
+We set the maximum simulation time to $120$s and conducted $10$ repeated trials for each scenario.
+A trial is considered successful if the robot achieves the goal within $120$s.
+Otherwise, the trial is considered a failure.
+Evaluation Metrics.
+The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), execution time (ET), and trajectory length under successful trials (TLS) to comprehensively assess the performance of different methods.
+SR refers to the portion of successful trials across 10 trials.
+OT denotes the average simulation time across all trials.
+If a trial fails, OT is recorded as the maximum simulation time, which is $120$s.
+OTS represents the average task completion time for successful trials.
+PT and ET indicate the average time spent on task planning and skill execution under successful completion, respectively.
+TLS refers to the average trajectory length of successful trials.
+SR, OT, OTS, and TLS indicate the overall effectiveness and efficiency of methods, while PT and ET reflect the computational efficiency and quality of task planning.
 Figure 5: Average success rate and overall time across all scenarios of different methods.
-Figure 5: Success cases of the proposed method in four interactive navigation tasks.
-The red phases indicate the robot interacts with objects, while the green phases represent the robot performing replanning based on new environmental observations.
-Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, Robotool and tree-planner achieve a higher success rate in interactive navigation tasks compared to traditional navigation methods.
-However, Robotool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks.
+Figure 6: Success cases of AINav in maze environments.
+The green markers above the figures indicate the robot interacts with objects, while the red markers represent the robot performing replanning based on new environmental observations.
+Figure 7: Success cases of AINav in Matterport3d, field ruin, and urban ruin.
+V-B Baseline Methods
+We compare the performance of AINav to the following baselines.
+Art-planner [wellhausen2023artplanner] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks.
+Hierarchical RL trains a high-level policy to function as the task planning module, which decides the low-level skill to use and the corresponding parameters to reach the goal.
+We also select two LLM-based method for comparison.
+The first is RoboTool [xu2023creative], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges.
+Since this method uses an open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output.
+The second method is Tree-planner [hu2023tree], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems.
+Because its plan sampling module requires global information, we include the initial object information as input.
+We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks.
+For a fair comparison, all methods use the same set of low-level skills as AINav.
+V-C Results and Analyses
+Simulation results are displayed in Table˜I.
+We can observe that AINav exhibits higher success rates and less overall time compared to other baselines in diverse scenarios.
+The Art-planner achieves high task success rates in low-difficulty scenarios of the maze environment with low execution time, showcasing its effective planning capabilities in traditional navigation tasks.
+However, in other scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks.
+Hierarchical RL can achieve limited success in low-difficulty scenarios, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
+Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, RoboTool and tree-planner achieve a higher success rate in interactive navigation tasks compared to Art-planner.
+However, RoboTool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks and disaster scenes.
 While tree-planner leverages a closed-loop design to adjust its decisions based on new observations, its performance is highly dependent on the quality of the sampled plans, resulting in instability in planning results.
@@
 Additionally, both methods require omniscient environmental information to generate task decomposition plans, rendering them ineffective in unknown environments.
-In contrast, our method not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information.
-This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios, especially in challenging situations that require replanning.
-Although node evaluation incurs additional planning time, the improved task planning quality and flexible replanning mechanism enable our method to achieve higher success rates and shorter overall completion time compared to other LLM-based approaches.
-As shown in Fig.
-4, AINav demonstrates the highest success rate and the lowest overall completion time across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
-We also present successful cases of our method across four tasks in Fig.
-5.
-In the box obstruction scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position.
-In the box usage and stair building scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which is unreachable for a quadruped robot under normal circumstances.
-In integrated task scenario, when confronted with various potential interaction tasks, the robot understands the current issue by VLM and intelligently interacts with the appropriate objects to facilitate task completion.
-Furthermore, we can observe that the robot is capable of responding to updated information, such as when the box fails to move or the robot identifies new objects and tasks.
-By leveraging an efficient closed-loop system, the proposed approach enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
-IV-D Ablation Studies
-Figure 6: Comparison results in ablation study.
-To further investigate the effectiveness of the proposed method, we conduct ablation study to compare AINav with the following four variants.
-1) AINav-Single, which only generates a single plan with LLM in the task planning module; 2) AINav-Skill, which directly generates skills without incorporating object-centric reasoning; 3) AINav-No, which removes the replanning module; 4) AINav-F, which only replans when a failure is encountered.
-We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building, which need interaction for task completion, to compare the success rate and overall time.
-The results of ablation studies are presented in Fig.
-6.
+In contrast, AINav not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information.
+This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios.
+Although node evaluation slightly increases planning time, the improved task planning quality and flexible replanning mechanism enable AINav to achieve higher success rates, shorter overall completion time, and trajectory length compared to all baseline approaches.
+As shown in Fig.˜5, AINav demonstrates the highest success rate and the lowest overall completion time among all approaches and across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
+We visualize the performance of AINav across different tasks and scenarios in Fig.˜6 and Fig.˜7.
+In the box obstruction and Matterport scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position.
+In the box usage, stair building, and ruin scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which are unreachable for a quadruped robot under normal circumstances.
+Notably, AINav’s advanced reasoning is highlighted in the stair-building and urban ruin tasks, where the robot must sequentially arrange boxes of varying heights in the correct order to climb to the prohibitively high platform.
+The successful completion of these tasks demonstrates the capacity of AINav for complex tool-use reasoning and execution.
+Furthermore, we can observe that the robot can respond to updated information, such as when the box fails to move or the robot identifies new objects and tasks.
+By leveraging an efficient closed-loop system, AINav enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
+TABLE II: Failure study of AINav in simulations
+Primary Category Sub-Category Rate
+Task Timeouts Scene understanding error $17.2\%$
+Incorrect task decomposition $34.5\%$
+Skill execution unfinished $24.1\%$
+Irrecoverable execution Intra-skill failure $10.3\%$
+Skill transition failure $13.8\%$
+We further conduct a failure study to analyze the current limitations of AINav.
+We categorize all failures observed during simulations into two primary types: task timeouts and irrecoverable execution failures, where task timeouts occur when the robot does not complete the task within the allocated time limit, and irrecoverable execution failures involve a critical error during primitive skill execution, such as the robot falling over.
+Table˜II summarizes the distribution of these failure modes across simulations.
+In summary, this failure study reveals that the primary bottlenecks for AINav lie in the efficiency of task planning and the seamless integration of low-level skills, highlighting clear directions for future improvement.
+Furthermore, the assumption of simplified visual processing and reliance on predefined constraints to determine skill executability limit AINav’s generalization in realistic scenarios, which will be the focus for future work to enhance robustness and adaptability.
+V-D Ablation Studies
+We conduct ablation studies in maze environments to compare AINav with the following four variants: 1) AINav-Single that only generates a single plan with LLM in the task planning module; 2) AINav-Skill that directly generates skills without incorporating object-centric reasoning; 3) AINav-No that removes the replanning module; 4) AINav-F that only replans when a failure is encountered.
+We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building in the maze environment, which need interaction for task completion, to compare the SR and OT.
+The results of ablation studies are presented in Fig.˜8.
 AINav-Single exhibits low success rates, as the single plan generated by the LLM is often flawed or inefficient due to inherent uncertainties of LLM, leading to task failures.
 Benefiting from the tree structure that enhances the robustness and planning quality, AINav-Skill achieves a reasonable success rate.
-However, it lacks a comprehensive understanding of object utilization in complex scenarios, leading to task decomposition plans inexecutable.
+However, it lacks a holistic understanding of object utilization in complex scenarios, leading to inexecutable task decomposition plans.
 AINav-No performs poorly because it lacks a replanning phase, making it unable to adapt to new observations.
-Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, often resulting in excessive time spent and task timeouts.
-By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving higher success rates and improved time efficiency.
-IV-E Performance of Primitive Skills
+Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, thus often resulting in excessive time spent and task timeouts.
+By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving the highest success rates and improved time efficiency.
+Figure 8: Comparison results in ablation study.
+VI Real-World Experiment
+We conduct real-world experiments to validate the performance of AINav in practical situations.
+We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection.
+We evaluate the effectiveness of AINav across five interactive navigation scenarios, as illustrated in Fig.˜10 (a).
+The first three scenarios are conducted in indoor environments: obstacle traversal, stair building, and corridor obstruction.
+In obstacle traversal, the robot is required to utilize a box as a stepping stone to surmount a 0.350.35m obstacle.
+The stair-building task requires the robot to arrange boxes of two different heights to construct a temporary staircase, enabling it to ascend a 0.500.50m platform.
+In corridor obstruction, two boxes obstruct a narrow passageway, one of which is heavy to push.
+The robot must proactively interact with these obstacles to create a navigable path and replan its strategy upon encountering the fixed box.
+The final two scenarios are set in outdoor environments, where the robot needs to utilize a box to traverse a 0.450.45m obstacle on the concrete floor or grass.
+We assume the quadruped’s maximum climbing height is 0.300.30m.
+We benchmark our method against the RoboTool baseline, which performs reasonably well in our simulations and has been validated through real-world experiments [xu2023creative].
+We adopt the SLAM Toolbox package to obtain the robot’s real-time pose.
+Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain object pose and dimension, thereby eliminating the risk of confounding factor detection errors.
+Figure 9: Box-pushing policy in different configurations.
+Each row demonstrates the pushing strategy under different initial box positions: left-shifted (top), centered (middle), and right-shifted (bottom).
+Figure 10: Visualization of real-world experiments.
+(a) shows various scenarios in real-world experiments, where green stars indicate the navigation goal.
+Highlighted boxes indicate movable objects, with green markers denoting light movable boxes and red markers denoting heavy boxes.
+(b)-(f) illustrate case studies for each scenario, in which the leftmost subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
+TABLE III: Metrics comparison in real-world experiments
+Scenario RoboTool AINav
+SR PT (s) ET (s) TLS (m) SR PT (s) ET (s) TLS (m)
+Obstacle traversal $20\%$ 16.4 40.2 6.8 80% 11.8 38.4 8.1
+Stair building $0\%$ - - - 20% 16.5 71.6 9.3
+Corridor obstruction $40\%$ 15.9 58.4 6.9 80% 18.1 53.7 6.4
+Urban climbing $60\%$ 14.3 45.2 7.0 80% 8.4 37.6 6.6
+Field climbing 60% 19.2 62.4 12.3 60% 11.3 60.4 9.4
+Average success rate $36\%$ 64%
+Fig.˜10 (b) showcases AINav’s capacity for dynamic replanning and creative tool utilization.
+Initially, the robot attempts a direct pushing strategy to withdraw a box blocking its path.
+Upon discovering that the box (red highlighted in Fig.˜10 (a)) is difficult to push, AINav replans and pushes a nearby smaller box as a support structure to scale the obstacle.
+This adaptation from unpredictable environmental information underscores AINav’s ability to traverse impassable barriers through tool use and robust strategic replanning.
+Fig.˜10 (c) showcases AINav’s advanced tool utilization capability.
+The robot constructs a makeshift staircase from boxes of varying heights to ascend a 0.5m platform, which is twice the height of the robot itself, showing AINav’s long-horizon reasoning and planning abilities.
+Although the task planner consistently generates viable solutions, the relatively low success rate is primarily attributed to execution failures, particularly during the challenging task of sequential climbing on non-fixed boxes.
+Addressing this execution limitation will be a key focus of our future work.
+Fig.˜10 (d) demonstrates the experiment on corridor obstruction.
+After an initial attempt to push one of two obstructive boxes fails, the robot strategically interacts with the second box, clearing a navigable path to its destination.
+Fig.˜10 (e) and (f) show the robot completing the interactive navigation task in the outdoor scenarios, employing task decomposition and skill execution to achieve the navigation goal, overcoming challenges that traditional navigation methods could not address.
+Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction.
+These real-world experiments demonstrate AINav’s capability to perform interactive navigation tasks across diverse scenarios, highlighting its proficiency in both complex tool utilization and dynamic replanning in response to new environmental information.
+Further details are available in our supplementary video.
+Table˜III presents the metric comparison across different scenarios in $5$ repeated experiments, where AINav achieves a higher success rate, less planning time, and execution time, highlighting the effectiveness and adaptivity of AINav in various tasks and both indoor and outdoor settings.
+To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig.˜9.
+We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement.
+This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
+VII Conclusion
+We introduce an LLM-based adaptive interactive navigation approach, AINav, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals.
+Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning.
+The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained.
+Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal.
+Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations.
+AINav has the potential to enhance robotic navigation capabilities, offering an effective and efficient solution for interactive navigation in challenging environments.
+Future work will work on tighter integration between foundation models and versatile motion capabilities for real-time embodied planning and control.
+We plan to leverage visual foundation models to identify task-related objects and extract corresponding properties, broadening the practical applicability of AINav to real-world scenarios.
+Moreover, we will expand the variety of primitive skills and enhance their robustness and interior seamless transition, ultimately aiming to extend our framework to a diverse range of robotic platforms and complex, long-horizon tasks.
+-A Prompts
+-A1 Object-Level Proposer
+You are a quadruped robot on the ground in a 3D world.
+Your goal is to navigate to a specific point in the 3D space.
+Your navigation goal is [goal point], and your current scene understanding is [scene understanding].
+There are several objects in the scene that you may utilize.
+We use two parameters, position and size, to represent the location and size of an object, respectively.
+Each object’s position is represented by a 3D vector [x, y, z].
+Each object’s size is represented as a 3D vector [length, width, height].
+[object description].
+You have a skill library containing the following skills and corresponding parameters: [skill library].
+You must follow these constraints: [constraints].
+Give me five different abstract plans for using objects to help you complete navigation tasks.
+The plan must include which objects you need to use, the sequence you use the objects, how to use the objects.
+You must analyze the problem step by step and show the thinking process.
+You must follow the following answer template:
+[begin of plan]
+Plan1: I need to use [objects]. First, … Second …
+Plan2: …
+[end of plan]
+-A2 Skill-Level Proposer
+You are a quadruped robot on the ground.
+Your goal is to navigate to a specific point in the 3D space.
+Your navigation goal is [goal point].
+You have a skill library containing the following skills and corresponding parameters: [skill library].
+Here are some abstract plans: [object-level plans], you need to generate detailed plans according to each abstract plan.
+Each step in the detailed plan consists of a skill with an abstract position like ’walk-to(’abstract position’)’.
+You must follow the following answer template:
+[begin of plan]
+Plan1: [
+(’step1’,’<skill>’),
+[end of plan]
+-A3 Evaluator
+You are a quadruped robot on the ground in a 3D world.
+Your goal is to navigate to a specific point in the 3D space.
+Your navigation goal is [goal point], and your current scene understanding is [scene understanding] There are several objects in the scene that you may utilize: [object description].
+You have a skill library containing the following skills and corresponding parameters: [skill library].
+Here are some detailed plans to accomplish the navigation task: … You need to evaluate the reward of each step in the plan based on its contribution to task completion, assigning a value between 0 and 1.
+Specifically, this involves simulating whether future steps can reach the goal under the current step to assess its impact.
+If the step does not satisfy the given constraints, the reward is 0.
+Constraints you must follow: [constraints].
+At the end of your response, reply to me with the following answer template:
+[begin of evaluation]
+Plan1: [
+(’step1’, ’<skill>’, ’reward: …’),
+[end of evaluation]
+-A4 Parameter Calculation
+The current plan is…, you need to calculate the 3D coordinates [x, y, z] of the abstract position in the plan based on the robot pose and the updated object information.
+The robot pose is [robot pose].
+The object information is [object description].
+You need to consider the spatial relationship between objects to obtain a reasonable position.
+You must calculate the position along each dimension step by step.
+-A5 Advisor
+You are a quadruped robot on the ground in a 3D world.
+Your goal is to navigate to a specific point in the 3D space.
+Your navigation goal is [goal point].
+There are several objects in the scene that you may utilize: [object description].
+Your current plan is: … Now, you have a new observation interpretation of the environment: … You need to determine whether to replan to modify your current plan based on the current plan, this new environmental observation, and the following criteria.
+You must analyze whether to reply, provide me with the reason, and respond with "Yes" or "No".
+Replanning criteria:
+* If the new observation is an execution failure of the current plan, then replanning is necessary.
+* If the new observation is a new object, you need to evaluate how this new object might help complete your task.
+If using it results in a more effective plan than your current one, you need to replan.
+* If the new observation is a revaluation of a previously known object, you need to determine whether this new information impacts your current plan.
+For example, if your plan requires climbing onto a box, but new observations show that the box is too high to climb, you need to replan.
+-A6 Arborist
+You are a quadruped robot on the ground in a 3D world.
+Your goal is to navigate to a specific point in the 3D space.
+Your navigation goal is [goal point].
+There are several objects in the scene that you may utilize: [object description].
+Constraints you must follow: [constraints].
+Your current plan is: … Now, you have a new observation of the environment: ….
+Here are replanning suggestions by the advisor: … You must make adjustments to your current plans based on the suggestions provided by the advisor, such as expanding new skills for new objects or pruning infeasible skills.
+-B Reward Design
+[…]
+TABLE IV: Definition of symbols.
+Symbol Description
+$\mathbf{q}_{j}$ Joint positions
+$\dot{\mathbf{q}_{j}}$ Joint velocities
+$\ddot{\mathbf{q}_{j}}$ Joint accelerations
+$\mathbf{q}^{*}_{j}$} Target joint positions
+$\mathbf{q}^{*,pre}_{j}$} Last target joint positions
+$\boldsymbol{\tau}_{j}$ Joint torques
+$\mathbf{v}_{b}$ Base linear velocity
+$\boldsymbol{\omega}_{b}$ Base angular velocity
+$\boldsymbol{x}_{b}$ Base position
+$\boldsymbol{\theta}_{b}$ Base heading
+$\boldsymbol{x}_{o}$ Object pose
+$\mathbf{v}^{*}_{b}$ Commanded base linear velocity
+$\boldsymbol{\omega}^{*}_{b}$ Commanded base angular velocity
+$\boldsymbol{x}^{*}_{b}$ Commanded base position
+$\boldsymbol{x}^{*}_{o}$ Commanded object pose
+$n_{c}$ Number of collisions
+$\mathbf{t}_{air,f}$ Feet air time
+TABLE V: Reward terms in walking skill.
+Reward terms Definition Weight
+Linear velocity tracking $\phi(\mathbf{v}^{*}_{b,xy}-\mathbf{v}_{b,xy})$ 11
+Angular velocity tracking $\phi(\boldsymbol{\omega}^{*}_{b,z}-\boldsymbol{\omega}_{b,z})$ $0.5$
+Linear velocity penalty $\mathbf{v}_{b,z}^{2}$ 22
+Angular velocity penalty $\boldsymbol{\omega}_{b,xy}||^{2}$ $0.05$
+Joint torques $\boldsymbol{\tau}_{j}||^{2}$ $0.00001$
+Joint accelerations $\ddot{\mathbf{q}_{j}}||^{2}$ $2.5*10^{-7}$
+Action rate $\mathbf{q}^{*}_{j}-\mathbf{q}^{*,pre}_{j}||^{2}$ $0.01$
+Collisions $n_{collision}$ 11
+Feet air time $\sum_{f=0}^{4}(\mathbf{t}_{air,f}-0.5)$ $0.125$
+TABLE VI: Reward terms in climbing skill.
+Reward terms Definition Weight
+[…]
+Move direction $\text{cos}\langle\mathbf{v}_{b},\boldsymbol{x}^{*}_{b}-\boldsymbol{x}_{b}\rangle$ 22
+Linear velocity penalty $\mathbf{v}_{b,z}^{2}$ 22
+Angular velocity penalty $\boldsymbol{\omega}_{b,xy}||^{2}$ $0.05$
+Joint torques $\boldsymbol{\tau}_{j}||^{2}$ $0.00001$
+Joint accelerations $\ddot{\mathbf{q}_{j}}||^{2}$ $2.5*10^{-7}$
+Action rate $\mathbf{q}^{*}_{j}-\mathbf{q}^{*,pre}_{j}||^{2}$ $0.01$
+Collisions $n_{collision}$ 11
+Feet air time $\sum_{f=0}^{4}(\mathbf{t}_{air,f}-0.5)$ $0.125$
+TABLE VII: Reward terms in navigation skill.
+Reward terms Definition Weight
+[…]
+Negative x-velocity penalty $\max(\mathbf{v}_{b,x},0)$ 22
+Collisions $n_{collision}$ $10$
+TABLE VIII: Reward terms in pushing skill.
+Reward terms Definition Weight
+[…]
+Object heading tracking $\boldsymbol{x}^{*}_{o,h}-\boldsymbol{x}_{o,h}|| 11$
+Negative x-velocity penalty $\max(\mathbf{v}_{b,x},0)$ 11
+Face to object $\text{cos}\langle\mathbf{\theta}_{b},\boldsymbol{x}_{o}-\boldsymbol{x}_{b}\rangle$ 22
+-C Performance of Primitive Skills
 (a) Climbing up/down skill
@@
 (a) and (b) show the climbing and navigation skills adapted to various terrains, while (c) and (d) showcase robust interaction capabilities with boxes of different shapes and sizes.
-We visualize the performance of various skills in Fig.
-7.
-In Fig.
-7(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks.
-Fig.
-7(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment.
-In Fig.
-7(c) and Fig.
-7(d), we showcase the robot pushes boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy.
-The box-pushing process typically consists of three steps: first, the robot approaches the box and positions itself so that the box is between the robot and the target position, preparing for the subsequent pushing action; second, the robot pushes the box toward the target point; and finally, the robot fine-tunes the box’s position to complete the task.
+We visualize the performance of various skills in Fig.˜11.
+In Fig.˜11(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks.
+Fig.˜11(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment.
+In Fig.˜11(c) and Fig.˜11(d), we showcase the robot pushing boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy.
 These versatile skills collectively form a comprehensive skill library that enables the robot to handle various sub-tasks in complex navigation scenarios.
-V Experiment
-Figure 8: Box-pushing policy in different configurations.
-(a) Different scenarios for real-world experiments.
-(b) Box obstruction
-(c) Box utilization: indoor
-(d) Box utilization: outdoor
-Figure 9: Visualization of real-world experiments.
-(a) shows various scenarios in real-world experiments.
-(b), (c) and (d) illustrate case studies for each scenario, in which the top-left subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
-We also conduct real-world experiments to validate the performance of AINav in practical situations.
-We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection.
-We evaluate the effectiveness of our method across three interactive navigation scenarios, as illustrated in Fig.
-9(a).
-The first scenario involves an indoor obstacle obstruction setting, simulating a navigation task in a narrow corridor blocked by movable obstacles, where the robot needs to remove the box to reach the target.
-The second scenario, also set in a confined indoor corridor, featured a fixed $0.35$m-height platform obstructing the path.
-Since we assume the maximum climbing height for the robot is $0.3$ meters, the robot needs to reason how to traverse heights that exceed this capability limit by interacting with objects.
-The third scenario is box utilization in an outdoor environment, setting with fixed obstacles blocking the path and $0.45$m height, presenting a more difficult challenge with a different layout.
-Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain accurate object descriptions, thereby mitigating the risk of task failure due to perception errors.
-Fig.
-9(b) demonstrates the experiment on obstacle obstruction.
-Initially, the robot detects a blocking box and navigates around it using local point cloud information and its navigation policy to approach the box.
-Subsequently, a pushing skill is employed to move the obstacle aside, creating a path to the target.
-Fig.
-9(c) illustrates how the robot uses a box to overcome a higher obstacle.
-After visually detecting the box, the task planning module decides to push the box next to the high platform as a step, enabling the robot to climb over the challenging obstacle and reach the target.
-Fig.
-9(d) shows the robot completing the interactive navigation task in the outdoor scenario, employing task decomposition and skill execution to achieve the target, overcoming challenges that traditional navigation methods could not address.
-Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction.
-These above real-world experiments demonstrate our method’s capability to complete navigation tasks in challenging environments.
-Further details are available in our supplementary video.
-We conduct $5$ repeated experiments in each scenarios, and the success rates are $100\%$ ($/5$ 5), $80\%$ ($4/5$), and $80\%$ ($4/5$), respectively, highlighting the robustness and superiority of AINav in various interactive tasks and both indoor and outdoor settings.
-To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig.
-8.
-We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement.
-This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
-VI Conclusion
-We introduce an LLM-based adaptive interactive navigation approach using a quadruped robot, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals.
-Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning.
-The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained.
-Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal.
-Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations.
-This instructive framework has the potential to enhance robotic navigation capabilities, offering an effective and computationally efficient solution for interactive navigation in challenging environments.
-Future directions encompass developing more generalizable interaction skills to accommodate broader tasks and scenarios, and achieving tighter integration between LLM/VLM and versatile control policies for more intelligent embodied planning and control.
+-D Case Visualization with Concrete Prompts
+We provide a detailed process for the obstacle traversal scenario in real-world experiments, including prompts and outputs for all LLM-based modules.
+This covers the initial task decomposition and the replanning stage in case of failure.
+Due to space limitations, we only present the most informative content in prompt and LLM output.
+Figure 12: Visualization of prompts and output of each module.